OpenStudy (anonymous):

What are the possible lengths and widths of a rectangle if its length is 8 centimeters longer than the width, and the area is less than or equal to 33 square centimeters?

OpenStudy (cwrw238):



OpenStudy (cwrw238):

this leads to the inequality x(x + 8) >= 33

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

so wait that is the answer

OpenStudy (cwrw238):

no you nee to solve this inequality x(x + 8) >= 33 x^2 + 8x - 33 >= 0 now factor the left hand side can you do that?

OpenStudy (anonymous):

no i need help